OpenStudy (anonymous):

Give an example of a rational function that has a horizontal asymptote at y = 0 and a vertical asymptote at, x = 2 and x = 1.

OpenStudy (anonymous):

so far i got f(x)=g(x)/f(x) so far i got f(x)=g(x)/(x-2)(x-1)